# A Comprehensive Survey on Instruction Following

### v7

@@ 7.2.2 Instruction consistency
This factor considers the instructions across the training tasks and test tasks. Keeping the instruction paradigm (e.g., abstractiveness) consistent is crucial in instruction following. Wei et al. (2022a) first investigated the performance impact of changing the instruction paradigm. They found that LLMs tuned on short instructions (i.e., task names) cannot generalize to longer sentence-style instructions (short ⇏normal-$\not\Rightarrow$ long). Similarly, Gu et al. (2022) observed the performance dropping when changing paragraph-style instructions to shorter sentence-style instructions at the test phase (long ⇏normal-$\not\Rightarrow$ short), further indicating the importance of instruction consistency.
Besides discrete instruction, maintaining the instruction paradigm is also critical for soft instruction, i.e., keeping the same-size prefix embedding when testing on unseen tasks Xu et al. (2022b). Interestingly, similar results were also found in the few-shot demonstrations (i.e., in-context learning), where the combination of input-output pairs or the number of demonstrations cannot be changed during training and evaluation Min et al. (2022a, b); Iyer et al. (2022). These phenomena raise a concern: although instruction-tuned LLMs are robust to tiny perturbations of instructions, they are vulnerable when facing more significant alterations, which is far behind human-level generalization.
@@ 7.2.3 Instruction diversity
Nevertheless, manually crafting instructions with diversity is expensive and usually hard to achieve due to the human annotation bias Huynh et al. (2021); Parmar et al. (2022). Owing to the excellent annotation quality of LLMs He et al. (2023); Pan et al. (2023), a considerable number of works began to employ models to compose innovative instructions Zhang et al. (2020, 2021); Honovich et al. (2022b). Although the model-generated instructions have been proven to contain more noise, benefiting from the diverse syntax structures Kitaev and Klein (2018), these instructions could still show complementary effects with the human-written instructions Wang et al. (2022c). More interestingly, Lou et al. (2023) proposed a new instruction-following dataset paradigm, where they employed LLMs to synthesize diverse task instructions for each input. Benefiting from this paradigm, the tuned LMs were forced to focus more on the instruction than the task input, achieving promising instruction-following performance. All of these results may imply the profitability of instruction diversity, even at the expense of the correctness of instructions.
@@ 7.2.4 Add demonstrations or not
Demonstrations, i.e., a couple of input-output examples, have been shown to be critical for the expressiveness of task instructions. For example, existing works found that adding a few positive demonstrations in the textual instructions could result in a significant performance improvement on the unseen tasks Yin et al. (2022); Deb et al. (2022), especially for the tasks occupying complex output space Mishra et al. (2022b); Wang et al. (2022d). Surprisingly, Gu et al. (2022) further found that models highly relied on few-shot demonstrations and even abandoned other useful resources (e.g., detailed task definition) when demonstrations were available. This prominence is perhaps because the LLMs prefer to exploit the more superficial patterns of the demonstrations rather than the other complex textual expressions Min et al. (2022b). In other words, at present, a comprehensive framework for accurately encoding pure instructions in the absence of demonstrations or task scaling remains elusive Lou and Yin (2023).
@@ 7.3.4 Emphasizing input-output mappings
For in-context learning, the model usually can not directly "learn" the input-output mapping from the given examples because there is no parameter update for the models. Therefore, one issue of in-context learning is that, when conducting instruction following, the demonstrations are not necessarily needed for the model to solve the task (i.e., even without the few-shot demonstrations, the model can still make predictions). Min et al. (2022b) also found that the model is more likely to “copy” the output candidate from the demonstrations, instead of truly learning the underlying mapping.
@@ 7.5 Data-wise Factor: Task Scale
The task scale often refers to the number of different training task categories in the dataset. Since “data-wise factor” also includes the scale of training instances, Wang et al. (2022d) investigated the impact of both task and instance scales. They found that instance scale (fixed task number, increasing the number of instances per task) can only bring a limited performance boost, while task scale is the key factor for instruction following, in line with the observations of other works Wei et al. (2022a); Chung et al. (2022). As illustrated in Figure 4, the same-size model with more tuning tasks usually gains better performance. However, the performance improvement of scaling up tasks is unstable, especially when the model size is too small (e.g., 0.08B Flan-T5). This phenomenon aligns with the discussion in § 7.1, we can draw a similar conclusion here: the profits of the task scale are highly governed by the model scale.
@@ 7.6 Main Takeaway: Dual-Track Scaling
Among all the factors discussed in this section, scaling is arguably the core factor that leads to the success of instruction following. Prior to LLM-based instruction following, scaling was mainly for deep learning models: from single-layer neural nets to multi-layer perceptions, from convolutional/recurrent neural networks to deep-layer transformers Hochreiter and Schmidhuber (1997); LeCun et al. (1998); Vaswani et al. (2017); Devlin et al. (2018). Along with the pretraining of massive raw text data, the ever-increasing models are expected to have encoded a vast amount of generic-purpose knowledge Zhou et al. (2023a). In the era of instruction following, where the community is more interested in cross-task generalization, merely scaling LLMs seems not enough. Thus, researchers take a parallel scaling: to collect more and more training tasks and labeled examples for each. We interpret this as a dual-track scaling. Overall, this dual-track scaling jointly seeks supervision to solve new tasks—the supervision either comes from LLMs’ pretraining or substantial training tasks. Despite its progress, some notable challenges remain in this area, which we will discuss in the next section.
@@ 8.2 Adversarial Instruction Attacks
Since instruction-tuned LLMs have been applied to various real-world scenarios, such as web agents and search engines Deng et al. (2023); Xie et al. (2023), the safety of LLMs’ generations is becoming more urgent. Simply conducting preference alignment or content filtering seems to be insufficient, especially for those super-strong LLMs. Thus, developing efficient defence methods is necessary for the current instruction-tuned models. Meanwhile, further deep analyses of LLMs’ vulnerability are also critical, potentially providing more insights into the defence.
@@ 9 Instruction-related Applications
In addition to the main body of our paper, we also survey some popular instruction-related application directions and put this section in the appendix to save space. If the final space allows, we will try to compress these discussions into the main body.
@@ 9.2 Data and Feature Augmentation
Task instructions are regarded as indirect supervision resources where sometimes superficial and assertive rules are embedded. These rules are also known as labeling functions that can be directly applied for annotations. [^6] For example, if “a very fair price” is sentiment-positive, every sentence with a similar adj-noun collocation as “fair price” will be positive as well. Therefore, some existing works also employed the instruction as a distant supervision to perform data or feature augmentation Srivastava et al. (2018); Hancock et al. (2018); Ye et al. (2020). For instance, Srivastava et al. (2017) used a semantic parser to convert natural language explanations into logical forms, and applied them on all instances in the dataset to generate additional binary features. While Wang et al. (2020) utilized the label explanations to annotate the raw corpus automatically and trained the classifier on the resulting noisy data.
@@ 9.3 Generalist Language Models
According to the definition of Artificial General Intelligence (AGI), the “generalist model” is usually a system that can be competent for different tasks and scalable in changeable contexts, which shall go far beyond the initial anticipations of its creators Wang and Goertzel (2007); Goertzel (2014). While specific to the NLP domain, a generalist language model is supposed to be an excellent multi-task assistant, that is skilled in handling a variety of real-world NLP tasks and different languages, in a completely zero/few-shot manner Arivazhagan et al. (2019); Pratap et al. (2020); Wei et al. (2022a). As numerous existing works demonstrated the incredible power of using instructions in cross-task generalization (Wei et al., 2022a; Sanh et al., 2022; Mishra et al., 2022b; Wang et al., 2022d; Chung et al., 2022, inter alia), the instruction is likely to become a breakthrough in achieving this ultimate goal.
Notably, the recent remarkable applications of instructions, namely InstructGPT, ChatGPT, and GPT-4 Ouyang et al. (2022); OpenAI (2022, 2023), also indicated a big step towards building generalist language models. However, unlike the other works that mainly employ instruction following, ChatGPT also adopts some other components, e.g., reinforcement learning with human feedback (RLHF). Although the answer to “which component contributes more to the dramatic results of ChatGPT” remains ambiguous and needs further investigation, we introduce some recent works highlighting the critical role of instruction following. For example, Chung et al. (2022) conducted extensive experiments to evaluate the human-preference alignments of PaLM Chowdhery et al. (2022). They found that, even without any human feedback, the instruction following significantly reduced the toxicity in the open-ended generations of PaLM, such as gender and occupation bias. In addition, some other works also solely employed creative instructions instead of human feedback and achieved notable cross-task results Bai et al. (2022b); Honovich et al. (2022a); Wang et al. (2022c).
@@ Appendix A. Instruction-following Datasets
UnifiedQA Khashabi et al. (2020) 05/2020 46 750 monolingual [Uncaptioned image] Human
CrossFit Ye et al. (2021) 04/2021 159 71,000 monolingual [Uncaptioned image] Human
Natural Instructions Mishra et al. (2022b) 04/2021 61 620 monolingual [Uncaptioned image] Human
Flan 2021 Wei et al. (2022a) 09/2021 62 4,400 monolingual [Uncaptioned image] Human
P3 Sanh et al. (2022) 10/2021 62 12,000 monolingual [Uncaptioned image] Human
MetaICL Min et al. (2022a) 10/2021 142 3,500 monolingual [Uncaptioned image] Human
ExMix Aribandi et al. (2022) 11/2021 107 500 monolingual [Uncaptioned image] Human
Super-Natural Instructions Wang et al. (2022d) 04/2022 1,613 5,000 multilingual [Uncaptioned image] Human
GLM Zeng et al. (2022) 10/2022 77 12,000 bilingual [Uncaptioned image] Human
Flan 2022 Longpre et al. (2023) 10/2022 1,836 15,000 multilingual [Uncaptioned image] Human
xP3 Muennighoff et al. (2022) 11/2022 71 81,000 multilingual [Uncaptioned image] Human
Unnatural Instructions Honovich et al. (2022a) 12/2022 117 64 monolingual [Uncaptioned image] InstructGPT
Self-Instruct Wang et al. (2022c) 12/2022 / 82 monolingual [Uncaptioned image] GPT-3
OPT-IML Iyer et al. (2022) 12/2022 2,207 18,000 multilingual [Uncaptioned image] Human
Alpaca Taori et al. (2023) 03/2023 / 52 monolingual [Uncaptioned image] InstructGPT
Baize Xu et al. (2023b) 04/2023 / 100 monolingual [Uncaptioned image] ChatGPT
Koala Geng et al. (2023) 04/2023 / / monolingual
@@
Alpaca-gpt4 Peng et al. (2023) 04/2023 / 113 bilingual [Uncaptioned image] GPT-4
Vicuna Chiang et al. (2023) 04/2023 / 76 monolingual
@@
Dolly Conover et al. (2023) 04/2023 / 15 monolingual [Uncaptioned image] Human
Oasst Köpf et al. (2023) 04/2023 / 84 multilingual [Uncaptioned image] Human
LongForm Köksal et al. (2023) 04/2023 / 27 monolingual
@@
Symbolic-Instruct Liu et al. (2023b) 04/2023 / 796 monolingual [Uncaptioned image] Human
LaMini Wu et al. (2023) 04/2023 / 2,580 monolingual [Uncaptioned image] ChatGPT
WizardLM Xu et al. (2023a) 04/2023 / 196 monolingual [Uncaptioned image] ChatGPT
COEDIT Raheja et al. (2023) 05/2023 / 82 monolingual [Uncaptioned image] Human
UltraChat Ding et al. (2023) 05/2023 / 1,500 monolingual [Uncaptioned image] ChatGPT
CoT Collection Kim et al. (2023) 05/2023 1,060 1,880 monolingual [Uncaptioned image] Codex
Dynosaur Yin et al. (2023) 05/2023 5,740 801 monolingual [Uncaptioned image] ChatGPT
MUFFIN Lou et al. (2023) 10/2023 / 68 monolingual
@@
Dynamics-of-Instruction Song et al. (2023) 10/2023 / 40 monolingual [Uncaptioned image] Human
CoachLM Liu et al. (2023d) 11/2023 / 2 monolingual [Uncaptioned image] Human
DEITA Liu et al. (2023c) 12/2023 / 10 monolingual [Uncaptioned image] ChatGPT
WaveCoder Yu et al. (2023b) 12/2023 4 20 monolingual

### v8

@@ 7.2.2 Instruction consistency
This factor considers the instructions across the training tasks and test tasks. Keeping the instruction paradigm (e.g., abstractiveness) consistent is crucial in instruction following. Wei et al. (2022a) first investigated the performance impact of changing the instruction paradigm. They found that LLMs tuned on short instructions (i.e., task names) cannot generalize to longer sentence-style instructions (short ⇏$\not\Rightarrow$ long). Similarly, Gu et al. (2023) observed the performance dropping when changing paragraph-style instructions to shorter sentence-style instructions at the test phase (long ⇏$\not\Rightarrow$ short), further indicating the importance of instruction consistency.
Besides discrete instruction, maintaining the instruction paradigm is also critical for soft instruction, i.e., keeping the same-size prefix embedding when testing on unseen tasks Xu et al. (2022a). Interestingly, similar results were also found in the few-shot demonstrations (i.e., in-context learning), where the combination of input-output pairs or the number of demonstrations cannot be changed during training and evaluation Min et al. (2022a, b); Iyer et al. (2022). These phenomena raise a concern: although instruction-tuned LLMs are robust to tiny perturbations of instructions, they are vulnerable when facing more significant alterations, which is far behind human-level generalization.
@@ 7.2.3 Instruction diversity
Nevertheless, manually crafting instructions with diversity is expensive and usually hard to achieve due to the human annotation bias Huynh et al. (2021); Parmar et al. (2023). Owing to the excellent annotation quality of LLMs He et al. (2023); Pan et al. (2023), a considerable number of works began to employ models to compose innovative instructions Zhang et al. (2020, 2021); Honovich et al. (2023b). Although the model-generated instructions have been proven to contain more noise, benefiting from the diverse syntax structures Kitaev and Klein (2018), these instructions could still show complementary effects with the human-written instructions Wang et al. (2023d). More interestingly, Lou et al. (2024) proposed a new instruction-following dataset paradigm, where they employed LLMs to synthesize diverse task instructions for each input. Benefiting from this paradigm, the tuned LMs were forced to focus more on the instruction than the task input, achieving promising instruction-following performance. All of these results may imply the profitability of instruction diversity, even at the expense of the correctness of instructions.
@@ 7.2.4 Add demonstrations or not
Demonstrations, i.e., a couple of input-output examples, have been shown to be critical for the expressiveness of task instructions. For example, existing works found that adding a few positive demonstrations in the textual instructions could result in a significant performance improvement on the unseen tasks Yin et al. (2022); Deb et al. (2022), especially for the tasks occupying complex output space Mishra et al. (2022b); Wang et al. (2022b). Surprisingly, Gu et al. (2023) further found that models highly relied on few-shot demonstrations and even abandoned other useful resources (e.g., detailed task definition) when demonstrations were available. This prominence is perhaps because the LLMs prefer to exploit the more superficial patterns of the demonstrations rather than the other complex textual expressions Min et al. (2022b). In other words, at present, a comprehensive framework for accurately encoding pure instructions in the absence of demonstrations or task scaling remains elusive Lou and Yin (2023).
@@ 7.3.4 Emphasizing input-output mappings
For in-context learning, the model usually cannot directly “learn” the input-output mapping from the given examples because there is no parameter update for the models. Therefore, one issue of in-context learning is that, when conducting instruction following, the demonstrations are not necessarily needed for the model to solve the task (i.e., even without the few-shot demonstrations, the model can still make predictions). Min et al. (2022b) also found that the model is more likely to “copy” the output candidate from the demonstrations, instead of truly learning the underlying mapping.
@@ 7.5 Data-wise Factor: Task Scale
The task scale often refers to the number of different training task categories in the dataset. Since “data-wise factor” also includes the scale of training instances, Wang et al. (2022b) investigated the impact of both task and instance scales. They found that instance scale (fixed task number, increasing the number of instances per task) can only bring a limited performance boost, while task scale is the key factor for instruction following, in line with the observations of other works Wei et al. (2022a); Chung et al. (2022). As illustrated in Figure 4, the same-size model with more tuning tasks usually gains better performance. However, the performance improvement of scaling up tasks is unstable, especially when the model size is too small (e.g., 0.08B Flan-T5). This phenomenon aligns with the discussion in § 7.1, we can draw a similar conclusion here: the profits of the task scale are highly governed by the model scale.
@@ 7.6 Main Takeaway: Dual-Track Scaling
Among all the factors discussed in this section, scaling is arguably the core factor that leads to the success of instruction following. Prior to LLM-based instruction following, scaling was mainly for deep learning models: from single-layer neural nets to multi-layer perceptions, from convolutional/recurrent neural networks to deep-layer transformers Hochreiter and Schmidhuber (1997); LeCun et al. (1998); Vaswani et al. (2017); Devlin et al. (2019b). Along with the pretraining of massive raw text data, the ever-increasing models are expected to have encoded a vast amount of generic-purpose knowledge Zhou et al. (2023a). In the era of instruction following, where the community is more interested in cross-task generalization, merely scaling LLMs seems not enough. Thus, researchers take a parallel scaling: to collect more and more training tasks and labeled examples for each. We interpret this as a dual-track scaling. Overall, this dual-track scaling jointly seeks supervision to solve new tasks—the supervision either comes from LLMs’ pretraining or substantial training tasks. Despite its progress, some notable challenges remain in this area, which we will discuss in the next section.
@@ 8 Challenges and Future Directions
8.1 The Tax of Instruction Alignment
The instruction following aims at taming the models to better assist humans in real-world tasks; therefore, in addition to pursuing ultimate performance, inference-time safety is also a crucial aspect for the instruction-tuned models (i.e., instruction alignment). Ouyang et al. (2022) defined “alignment” with three criteria — Helpful, Honest, and Harmless (HHH), which has been widely considered by the previous instruction tuning models and datasets Bai et al. (2022b); Yin et al. (2023); Wang et al. (2023d); Lou et al. (2024). However, alignment can also bring “tax” to the instruction-tuned models. For example, Bekbayev et al. (2023) found that the well-aligned answers provided in the instruction following datasets can drop the model’s performance a lot on various task benchmarks. This implies a trade-off between performance and safety for instruction following, which requires careful consideration.
@@ 8.3 Adversarial Instruction Attacks
Since instruction-tuned LLMs have been applied to various real-world scenarios, such as web agents and search engines Deng et al. (2023); Xie et al. (2024a), the safety of LLMs’ generations is becoming more urgent. Simply conducting preference alignment or content filtering seems to be insufficient, especially for those super-strong LLMs. Thus, developing efficient defence methods is necessary for the current instruction-tuned models. Meanwhile, further deep analyses of LLMs’ vulnerability are also critical, potentially providing more insights into the defence.
@@ 8.5 Learning to Follow Instruction rather than Merely Generating Y
8.6 Multi-Lingual Instruction Following
Intuitively, the instruction following is language-agnostic capacity for the language models, which means that it is also possible for the multi-lingual language models to follow the same semantic instructions with different languages. For example, Kew et al. (2023) found LLMs tuned with more than three languages exhibit stronger instruction following capacity, implying the benefits of multi-lingual instruction tuning. Unfortunately, most of the current open-sourced instruction following datasets and foundation models are English-centric (as shown in Table 6). Therefore, the release of high-quality multi-lingual instruction tuning datasets (with pair translation) should be valuable for future research, as also mentioned by Peng et al. (2023).
@@ 9 Instruction-related Applications
In addition to the main body of our paper, we also survey some popular instruction-related application directions to inspire future board-wide utilization for instruction following.
@@ 9.2 Data and Feature Augmentation
Task instructions are regarded as indirect supervision resources where sometimes superficial and assertive rules are embedded. These rules are also known as labeling functions that can be directly applied for annotations. [^6] For example, if “a very fair price” is sentiment-positive, every sentence with a similar adj-noun collocation as “fair price” will be positive as well. Therefore, some existing works also employed the instruction as a distant supervision to perform data or feature augmentation Srivastava et al. (2018); Hancock et al. (2018); Ye et al. (2020). For instance, Srivastava et al. (2017) used a semantic parser to convert natural language explanations into logical forms, and applied them on all instances in the dataset to generate additional binary features. Wang et al. (2020) utilized the label explanations to annotate the raw corpus automatically and trained the classifier on the resulting noisy data.
@@ 9.3 Generalist Language Models
According to the definition of Artificial General Intelligence (AGI), the “generalist model” is usually a system that can be competent for different tasks and scalable in changeable contexts, which shall go far beyond the initial anticipations of its creators Wang and Goertzel (2007); Goertzel (2014). While specific to the NLP domain, a generalist language model is supposed to be an excellent multi-task assistant, that is skilled in handling a variety of real-world NLP tasks and different languages, in a completely zero/few-shot manner Arivazhagan et al. (2019); Pratap et al. (2020); Wei et al. (2022a). As numerous existing works demonstrated the incredible power of using instructions in cross-task generalization (Wei et al., 2022a; Sanh et al., 2022; Mishra et al., 2022b; Wang et al., 2022b; Chung et al., 2022, inter alia), the instruction is likely to become a breakthrough in achieving this ultimate goal.
Notably, the recent remarkable applications of instructions, namely InstructGPT, ChatGPT, and GPT-4, also indicated a big step towards building generalist language models. For example, during the pretraining of LLama-2, Touvron et al. (2023) utilized the idea of context distilling to inculcate instructions within LLMs, thus addressing the inconsistency issue of instruction following in the multi-turn dialogue situation. OpenAI GPT-series adopt RLHF to align the model’s preference with human instructions, where feedback supervision plays a big role. Although the answer to “Is instruction or human feedback, that contributes more to the performance of ChatGPT?” remains ambiguous and needs further investigation, we introduce some recent works highlighting the critical role of instruction following. For example, Chung et al. (2022) conducted extensive experiments to evaluate the human-preference alignments of PaLM Chowdhery et al. (2023). They found that, even without any human feedback, the instruction following significantly reduced the toxicity in the open-ended generations of PaLM, such as gender and occupation bias. In addition, some other works also solely employed creative instructions instead of human feedback and achieved notable cross-task results Bai et al. (2022b); Honovich et al. (2023a); Wang et al. (2023d). Furthermore, as the knowledge conflict problem of LLMs has a significant impact on the applications of instruction-tuned models Xie et al. (2024a), in order to make the LLMs more generalist and useful in the real world, recent works also utilized the idea of the instruction following to enhance the retrieval-augmented language models, and vice versa, improve the instructions by adopting retrieved knowledge Lin et al. (2023).
@@ Appendix A. Instruction-following Datasets
UnifiedQA Khashabi et al. (2020) 05/2020 46 750 monolingualmonolingual [Uncaptioned image] Human
CrossFit Ye et al. (2021) 04/2021 159 71,000 monolingualmonolingual [Uncaptioned image] Human
Natural Instructions Mishra et al. (2022b) 04/2021 61 620 monolingualmonolingual [Uncaptioned image] Human
Flan 2021 Wei et al. (2022a) 09/2021 62 4,400 monolingualmonolingual [Uncaptioned image] Human
P3 Sanh et al. (2022) 10/2021 62 12,000 monolingualmonolingual [Uncaptioned image] Human
MetaICL Min et al. (2022a) 10/2021 142 3,500 monolingualmonolingual [Uncaptioned image] Human
ExMix Aribandi et al. (2022) 11/2021 107 500 monolingualmonolingual [Uncaptioned image] Human
Super-Natural Instructions Wang et al. (2022b) 04/2022 1,613 5,000 multilingualmultilingual [Uncaptioned image] Human
GLM Zeng et al. (2022) 10/2022 77 12,000 bilingualbilingual [Uncaptioned image] Human
Flan 2022 Longpre et al. (2023) 10/2022 1,836 15,000 multilingualmultilingual [Uncaptioned image] Human
xP3 Muennighoff et al. (2023) 11/2022 71 81,000 multilingualmultilingual [Uncaptioned image] Human
Unnatural Instructions Honovich et al. (2023a) 12/2022 117 64 monolingualmonolingual [Uncaptioned image] InstructGPT
Self-Instruct Wang et al. (2023d) 12/2022 / 82 monolingualmonolingual [Uncaptioned image] GPT-3
OPT-IML Iyer et al. (2022) 12/2022 2,207 18,000 multilingualmultilingual [Uncaptioned image] Human
Alpaca Taori et al. (2023) 03/2023 / 52 monolingualmonolingual [Uncaptioned image] InstructGPT
Baize Xu et al. (2023b) 04/2023 / 100 monolingualmonolingual [Uncaptioned image] ChatGPT
Koala Geng et al. (2023) 04/2023 / / monolingualmonolingual
@@
Alpaca-gpt4 Peng et al. (2023) 04/2023 / 113 bilingualbilingual [Uncaptioned image] GPT-4
Vicuna Chiang et al. (2023) 04/2023 / 76 monolingualmonolingual
@@
Dolly Conover et al. (2023) 04/2023 / 15 monolingualmonolingual [Uncaptioned image] Human
Oasst Köpf et al. (2023) 04/2023 / 84 multilingualmultilingual [Uncaptioned image] Human
LongForm Köksal et al. (2023) 04/2023 / 27 monolingualmonolingual
@@
Symbolic-Instruct Liu et al. (2023b) 04/2023 / 796 monolingualmonolingual [Uncaptioned image] Human
LaMini Wu et al. (2024) 04/2023 / 2,580 monolingualmonolingual [Uncaptioned image] ChatGPT
WizardLM Xu et al. (2023a) 04/2023 / 196 monolingualmonolingual [Uncaptioned image] ChatGPT
COEDIT Raheja et al. (2023) 05/2023 / 82 monolingualmonolingual [Uncaptioned image] Human
UltraChat Ding et al. (2023) 05/2023 / 1,500 monolingualmonolingual [Uncaptioned image] ChatGPT
CoT Collection Kim et al. (2023) 05/2023 1,060 1,880 monolingualmonolingual [Uncaptioned image] Codex
Dynosaur Yin et al. (2023) 05/2023 5,740 801 monolingualmonolingual [Uncaptioned image] ChatGPT
MUFFIN Lou et al. (2024) 10/2023 / 68 monolingualmonolingual
@@
Dynamics-of-Instruction Song et al. (2023) 10/2023 / 40 monolingualmonolingual [Uncaptioned image] Human
CoachLM Liu et al. (2023d) 11/2023 / 2 monolingualmonolingual [Uncaptioned image] Human
DEITA Liu et al. (2023c) 12/2023 / 10 monolingualmonolingual [Uncaptioned image] ChatGPT
WaveCoder Yu et al. (2023b) 12/2023 4 20 monolingualmonolingual